Determine
the centre of mass and the reduced mass of
the earth moon system.
1
Expert's answer
2018-06-24T10:56:08-0400
r= m2*R/(m1+m2)=0.0735*10^24 * 384.467*10^6/ (5.976*10^24+0.0735*10^24)=4672*10^3 metres

Answer: The center of mass is inside the earth at a distance 4672000 metres from its center.